OpenStudy (anonymous):

I'm really stressed out, could someone please explain this problem to me? C.) Graph the system of equations and show the point of intersection.

OpenStudy (anonymous):

wait, what is the problem?

OpenStudy (anonymous):

The session is timed out

OpenStudy (anonymous):

oh, um it needs a password

OpenStudy (bibby):

We're not in your school. Take a screenshot...

OpenStudy (anonymous):

sorry, I don't know the password

OpenStudy (anonymous):

ok let me take a screenshot

OpenStudy (anonymous):

A.) Complete the table of values: x 100 + 23x 90(1.2^x) 0 - 100, 90 1 - 123, 108 2 - 146, 129.6 3 - 169, 155.52 4 - 192, 186.6 5 - 215, 223.95 Graph the system of equations and show the point of intersection.

OpenStudy (anonymous):

@bibby

OpenStudy (bibby):

graph f(x) = 100 + 23x and g(x) = 90(1.2^x) to get the intersection you can just set them equal to each other IIRC

OpenStudy (anonymous):

How do I set them equal to each other?

OpenStudy (anonymous):

If you put them equal to each other you get .x is approximately equal to -1.20871 and 4.4334. I don't know if that's what you want or not.

OpenStudy (anonymous):

Thank you